OpenStudy (anonymous):

The area of a rectangular garden is 336 square meters. The width is 2 meters longer than 1/2 of the length. Find the length and the width.

OpenStudy (anonymous):

width=(1/2)x+2 where x is the length 336=x*((1/2)x+2) 336=(1/2)x^2+2 668=x^2 x=sqrt(668) I think?

OpenStudy (anonymous):

Sorry DiSc02, I need the length and width.

OpenStudy (anonymous):

Let L and w be the length and width respectively. \[\left\{L w=336,w=\frac{L}{2}+2\right\} \]\[\{L=24,w=14\} \]

OpenStudy (anonymous):

I solved it. The length is 24 and the width is 14

OpenStudy (anonymous):

lol thanks Robtobey I got the same answer

OpenStudy (anonymous):

let the length be x metres therefore \[width = \frac{1}{2}\times x +2\] Area of rectangular garden = length *width i.e. \[336 = x \times ( \frac{1}{2} x + 2)\] \[336 = x \times \frac{1}{2}(x + 4)\] \[336 \times 2 = x \times (x + 4)\] \[672 = x^2 + 4x \rightarrow x^2 + 4x -672 =0\] \[x^2 + 28x -24x -672 =0 \rightarrow x(x+28)-24(x+28) =0\] \[\rightarrow (x+28)(x-24)=0\] i.e. (x+28)= 0 or (x-24)=0 i.e. x= -28 or x= 24 but -28 cannnot be the length as it is a negative number Hence x= 24 is the length of the rectangular garden Therefore Width = \[Width = \frac{1}{2} \times 24 + 2 = 12 + 2 = 14 units\] Hence length is 24 units and width is 14 units @Tfreeland88

OpenStudy (anonymous):

Thank you for all your work dpasingh!

OpenStudy (anonymous):

@Tfreeland88 Medal please.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

just click on the best response button beside my name

OpenStudy (anonymous):

Oh, someone had it before you, is there a way to still reward you?